# Functional Clustering for Longitudinal Associations between Social Determinants of Health and Stroke Mortality in the US

### v2

@@ Abstract
Understanding longitudinally changing associations between Social determinants of health (SDOH) and stroke mortality is crucial for timely stroke management. Previous studies have revealed a significant regional disparity in the SDOH – stroke mortality associations. However, they do not develop data-driven methods based on these longitudinal associations for regional division in stroke control. To fill this gap, we propose a novel clustering method for SDOH – stroke mortality associations in the US counties. To enhance interpretability and statistical efficiency of the clustering outcomes, we introduce a new class of smoothness-sparsity pursued penalties for simultaneous clustering and variable selection in the longitudinal associations. As a result, we can identify important SDOH that contribute to longitudinal changes in the stroke mortality, facilitating clustering of US counties into several regions based on how these SDOH relate to stroke mortality. The effectiveness of our proposed method is demonstrated through extensive numerical studies. By applying our method to a county-level SDOH and stroke mortality longitudinal data, we identify 18 important SDOH for stroke mortality and divide the US counties into two clusters based on these selected SDOH. Our findings unveil complex regional heterogeneity in the longitudinal associations between SDOH and stroke mortality, providing valuable insights in region-specific SDOH adjustments for mitigating stroke mortality.
Keywords: Functional clustering, Mixture model, Variable selection, Regularized expectation-maximization algorithm, Stroke
@@ 1 Introduction
The burden of stroke in the United States is enormous. As one of the most prevalent cardiovascular diseases, stroke remains consistently among the top five causes of death in the country (Koton et al., 2014), leading to over 130,000 fatalities annually (Holloway et al., 2014). In the effort to prevent stroke deaths, Social Determinants of Health (SDOH) – encompassing economic, social, and environmental conditions where people live, learn, work, and play (Havranek et al., 2015) – have garnered significant attention due to their strong associations with stroke mortality (Powell-Wiley et al., 2022). Notably, in 2015, the impact of SDOH on stroke mortality was highlighted in an important scientific statement by the American Heart Association (AHA) (Mozaffarian et al., 2015). Furthermore, since 2019, the AHA has consistently emphasized the importance of SDOH in its Annual Heart Disease and Stroke Statistics Report, spanning across all chapters (Benjamin et al., 2019). These declarations underscore the urgency of understanding the associations between SDOH and stroke mortality. This understanding is crucial for informing targeted interventions aimed at controlling stroke mortality by addressing underlying social, economic, and environmental factors.
@@
Recent findings have revealed a clear regional disparity in the associations between SDOH and stroke mortality. For instance, Zelko et al. (2023) identified state-wise differences in these associations. Additionally, Villablanca et al. (2016) and Son et al. (2023) observed significant disparities in the associations between rural and urban areas. To address such regional disparities, various state-wise (Gebreab et al., 2015) and rural-urban strategies (Labarthe et al., 2014; Record et al., 2015; Kapral et al., 2019) have been developed for region–specific stroke management. However, the existing literature does not guarantee that areas within a divided region share similar associations between SDOH and stroke mortality. Hence, employing a uniform SDOH adjustment in divided regions may not be an effective strategy for preventing stroke disease. Moreover, many SDOH have been found to exhibit longitudinal changes in their associations with stroke mortality (He et al., 2021), a factor crucial for timely policymaking in stroke management. Targeting these longitudinal associations between SDOH and stroke mortality, clustering becomes a vital tool for determining reasonable regional divisions for stroke prevention. Nonetheless, this remains an unsolved issue requiring further investigation.
In this work, we propose a novel method for clustering longitudinal associations between SDOH and stroke mortality. In general, the clustering is implemented based on the similarity among associations from different counties in the US, and the resulting clustering outcomes can then be used to inform region-specific prevention measures for controlling stroke mortality.
@@
To achieve this, we utilize county-level longitudinal data of SDOH and stroke mortality in the US, and treat the longitudinally observed data in each county as functional data. As a result, the concurrent associations between SDOH and stroke mortality are inherently functional objects, which can be effectively modeled using functional regression models, where the coefficients are also functional, capturing the relationship between SDOH and stroke mortality over time. For the clustering process, we model the functional coefficients using a finite mixture model, leading to a finite mixture of functional regression models. This approach enables the clustering of counties in the US into different regions based on the relationship between their SDOH and stroke mortality.
@@
Our task essentially differs from the common clustering procedures for longitudinal data (Jacques and Preda, 2013, 2014; Liang et al., 2021), which primarily apply to observable longitudinal outcomes. In contrast, the county-level longitudinal associations between SDOH and stroke mortality in our case are unobservable. For clustering of these associations, one might consider applying the finite mixture of regression models (Jacobs et al., 1991; Jiang and Tanner, 1999) to the SDOH and stroke mortality longitudinal data. However, this method cannot capture longitudinal changes in the associations within the clustering procedures, potentially resulting in unreliable clustering outcomes due to the disregard of longitudinal signals. In light of functional data analysis (Ramsay and Silverman, 2002, 2005), some studies proposed a finite mixture of functional regression models for clustering longitudinal associations (Yao et al., 2011; Lu and Song, 2012). Nonetheless, they do not address the issue of collinearity among covariates in their functional regression models. In our study, the collection of SDOH serves as functional covariates and exhibits significant collinearity, stemming from their derivation from multiple domains (e.g., social, economic, and environmental domains). In this case, ignoring colinearity among the SDOH data may lead to misspecification in the functional regression model. This oversight could compromise the accuracy of the resultant clustering outcome.
@@
Furthermore, to enhance interpretability, existing studies analyzing associations between SDOH and stroke mortality often adopt a pre-selection step on the SDOH covariates (Tsao et al., 2022, 2023), as their number is usually large, containing hundreds of variables. In our case, the pre-selection of longitudinal SDOH data can be achieved through variable selection in functional regression models (Wang et al., 2008; Kong et al., 2016; Goldsmith and Schwartz, 2017). However, these methods generally assume that associations between covariates and responses are invariant across different samples. This assumption limits their direct applicability to our case, where SDOH covariates may exhibit distinct associations with stroke mortality among different counties. Moreover, performing selections of SDOH prior to clustering may introduce biases for the subsequent clustering outcome, as the selection of SDOH is unrelated to the clustering process. On the other hand, without proper selection of SDOH, the clustering process may be statistically inefficient due to the complex structure of longitudinal SDOH data, which possess high-dimensionality and functional nature simultaneously. To accommodate such a complicated structure, it would be beneficial to connect variable selections of SDOH to the clustering of longitudinal associations, yet this topic is rarely discussed in the literature.
@@
In this article, we introduce a novel method for simultaneous clustering and variable selection of longitudinal associations between SDOH and stroke mortality in US counties. Our method, based on a finite mixture of functional linear concurrent models (FMFLCM), incorporates a new class of smoothness-sparsity pursued penalties to address the functional nature and high-dimensionality of the SDOH data. These penalties are designed to borrow information from distinct clusters of the FMFLCM, thereby enhancing statistical efficiency for both clustering and variable selection, and simultaneously addressing the collinearity issue among the SDOH data. For the estimation, we develop a novel regularized expectation-maximization (REM) algorithm by incorporating the proposed penalties. This approach allows for the clustering of longitudinal associations while embedding a variable selection step of SDOH covariates. As a result, the cluster memberships and the selected SDOH can be iteratively updated within the REM algorithm. This iterative updating process helps mitigate potential biases stemming from the selected covariates during the clustering process. Through these procedures, we provide a novel data-driven method for county-level regional division, aiming to offer insights into region-specific stroke prevention measures for US counties.
@@
The remainder of this article is organized as follows. In Section 2, we begin by introducing the SDOH and stroke mortality dataset, and then proceed to demonstrate some of their data features in Section 2.1. Following this, we present the FMFLCM in Section 2.2, followed by the demonstration of sparsity and smoothness pursued penalties in Section 2.3, and the REM in Section 2.4. In Section 3, we conduct simulation studies to compare the proposed methods with some competing approaches in terms of clustering performance, variable selection, and parameter estimation. In Section 4, we apply the proposed clustering method to our dataset and present the clustering result, along with the estimation relating to the selected SDOH. Finally, We provide conclusions and discussion in Section 5.
@@ 2.1 Data Source
In 2020, the Agency for Healthcare Research and Quality (AHRQ) compiled and released a SDOH database for a better understanding of community-level factors, healthcare quality and delivery, and individual health. The SDOH database contains yearly records of $345$ SDOH collected from 3226 counties from 2009 to 2018. These SDOH are classified into 5 domains: (1) social context, such as age, race and ethnicity, and veteran status; (2) economic context, such as income and unemployment; (3) education; (4) physical infrastructure, such as housing, food insecurity, and transportation; and (5) health care contexts, such as health insurance coverage and health care access. To study the association between the SDOH and stroke mortality, we connect the SDOH database with a stroke mortality data provided by the Interactive Atlas of Heart Disease and Stroke at the Centers for Disease Control and Prevention (CDC) on the county level. The stroke mortality database was originally compiled from 2 data sources: (1) the National Vital Statistics System at the National Center for Health Statistics, and (2) the hospital discharge data from the Centers for Medicare & Medicaid Services’ Medicare Provider Analysis and Review (MEDPAR) file. All data and materials used in this analysis are publicly available at the AHRQ website: https://www.ahrq.gov/sdoh/index.html and CDC website https://www.cdc.gov/dhdsp/maps/atlas/index.htm. Since the AHRQ database is HIPAA (Health Insurance Portability and Accountability Act) compliant, our data do not require to be reviewed by an institutional review board.
It’s worth noting that the SDOH dataset contains missing values. Following the approach of a previous study on the dataset (Son et al., 2023), we exclude the SDOH variables that have a missing proportion of more than 60%. The average missing proportion of the remaining variables is 3.6%. For these missing values, we employ a k-nearest neighbors (KNN) method (Kowarik and Templ, 2016) to impute the SDOH data, ensuring the SDOH and stroke mortality data are aligned with time for each county. The detailed implementation of the KNN is provided in Part A of the Supplementary Material.
@@
Figure 1: The correlation matrix for the SDOH variables (indexed by numbers). The correlation of the longitudinal data between the $j$th SDOH and $j^{\prime}$th SDOH is calculated as $C_{j,j^{\prime}}/\sqrt{C_{j,j}C_{j^{\prime},j^{\prime}}}$, where $C_{j,j^{\prime}}=n^{-1}\sum_{i=1}^{n}S_{i}^{-1}\sum_{s=1}^{S_{i}}\big{\{}X_{ij}(t_{is})-\hat{\mu}_{j}(t_{is})\big{\}}\big{\{}X_{ij^{\prime}}(t_{is})-\hat{\mu}_{j^{\prime}}(t_{is})\big{\}}$ with $\hat{\mu}_{j}(\cdot)$ being the estimated mean function for the $j$th county. Here, $X_{ij}(t_{is})$ is the observed data of the $j$th SDOH at the $i$th county on year $t_{is}$, for $i=1,\dots,n$, $j=1,\dots,p$, and $s=1,\dots,S_{i}$. The definition of $X_{ij}(t_{is})$, $n$, $p$, and $S_{i}$ can be found in Section 2.2.
As mentioned previously, significant collinearity may exist among the SDOH data. This phenomenon is observed in our dataset, where the longitudinal data between some of the SDOH exhibit significant correlations as illustrated in Figure 1. Besides, coinciding with the previous studies (Villablanca et al., 2016; He et al., 2021; Son et al., 2023), evident regional disparities and longitudinal variations can also observed in the SDOH – stroke mortality associations from our dataset. To demonstrate this, we calculate the yearly Pearson correlation between stroke mortality and the percentage of Asian and Pacific Island language speakers (PAPL), one principal sociocultural factor in the SDOH. To calculate the yearly correlations, we first divide the counties in the US into several regions, and presume that the counties in a divided region exhibit the same SDOH – stroke mortality correlation. We consider two types of divisions: state-wise division (Zelko et al., 2023) and rural-urban division (Son et al., 2023). The yearly correlations from 2009-2018 for each state, or for rural and urban areas, are presented in panels (B) and (D) of Figure 2, respectively. Besides, we also present the geographical maps of correlations for the years 2009, 2013, and 2018 in panels (A) and (C) of Figure 2, respectively for two types of division strategies.
@@
Figure 2: State-wise and rural-urban Pearson correlations between the stroke mortality and PAPL. Panels (A) and (C) are the geographical maps of correlations for the years 2009, 2013, and 2018, respectively for two types of division strategies. Panels (B) and (D) are the state-wise and rural-urban longitudinal correlations from 2009 – 2018, respectively.
@@
From the maps in Figure 2, we observe significant regional disparities in the correlations among different states, as well as between rural and urban areas. These correlations all exhibit changes over time. In addition, we find that the longitudinal correlations in rural and urban areas (shown in panel (D) of Figure 2) are consistently negative over time, while the correlations from some states (such as Nevada, as seen in panel (B) of Figure 2) are mostly positive. These results suggest that the outcomes of longitudinal associations are highly sensitive to the region division strategy, and different region divisions may lead to inconsistent conclusions for the stroke management. As such, we require a reasonable data-driven method for the region division in exploring SDOH – stroke mortality associations.

### v6

@@ Abstract
Understanding the longitudinally changing associations between Social Determinants of Health (SDOH) and stroke mortality is essential for effective stroke management. Previous studies have uncovered significant regional disparities in the associations between SDOH and stroke mortality. However, existing studies have not utilized longitudinal associations to develop data-driven methods for regional division in stroke control. To fill this gap, we propose a novel clustering method to analyze SDOH-stroke mortality associations across U.S. counties. To enhance the interpretability of the clustering outcomes, we introduce a novel regularized expectation-maximization algorithm equipped with sparsity- and smoothness-pursued penalties, aiming at simultaneous clustering and variable selection in longitudinal associations. As a result, we can identify key SDOH that contributes to longitudinal changes in stroke mortality. This facilitates the clustering of U.S. counties based on the associations between these SDOH and stroke mortality. The effectiveness of our proposed method is demonstrated through extensive numerical studies. By applying our method to longitudinal data on SDOH and stroke mortality at the county level, we identify 18 important SDOH for stroke mortality and divide the U.S. counties into two clusters based on these selected SDOH. Our findings unveil complex regional heterogeneity in the longitudinal associations between SDOH and stroke mortality, providing valuable insights into region-specific SDOH adjustments for mitigating stroke mortality.
Keywords: Model-based clustering, Functional regression model, Variable selection, Regularized expectation-maximization algorithm, Stroke mortality
@@ 1 Introduction
The burden of stroke in the United States is enormous. As one of the most prevalent cardiovascular diseases, stroke remains consistently among the top five causes of death in the country (Koton et al. (2014)), leading to over 130,000 fatalities annually (Holloway et al. (2014)). In the effort to prevent stroke deaths, Social Determinants of Health (SDOH)—encompassing economic, social, and environmental conditions where people live, learn, work, and play (Havranek et al. (2015))—have garnered significant attention due to their strong associations with stroke mortality (Reshetnyak et al. (2020); Skolarus et al. (2020); Powell-Wiley et al. (2022)). In 2013, Centers for Disease Control and Prevention (CDC) released an important report on health disparities (Frieden et al. (2013)), where the relationships between SDOH and stroke mortality were extensively documented. In 2015, the impact of SDOH on stroke mortality was highlighted in an important scientific statement by the American Heart Association (AHA) (Mozaffarian et al. (2015)). Since 2019, the AHA has consistently emphasized the importance of SDOH in its Annual Heart Disease and Stroke Statistics Report, spanning across all chapters (Benjamin et al. (2019)). These declarations underscore the urgency of understanding the associations between SDOH and stroke mortality. This understanding is crucial for informing targeted interventions aimed at controlling stroke mortality by addressing underlying social, economic, and environmental factors.
@@
Recent findings have revealed a clear regional disparity in the associations between SDOH and stroke mortality. For instance, Zelko et al. (2023) identified statewise differences in these associations. Additionally, Villablanca et al. (2016) and Son et al. (2023) observed significant disparities in the associations between rural and urban areas. To address such regional disparities, various statewise (Gebreab et al. (2015)) and rural-urban strategies (Labarthe et al. (2014); Record et al. (2015); Kapral et al. (2019)) have been developed for region-specific stroke management. However, the existing literature primarily adopts prespecified regional division strategies when analyzing regional disparities, which do not guarantee that areas within a divided region share similar associations between SDOH and stroke mortality. This can potentially lead to ineffective regional division strategies for stroke control. Moreover, many SDOH exhibit longitudinal changes in their associations with stroke mortality (He et al. (2021); Imoisili (2024)), a factor crucial for timely policymaking in stroke management. Targeting these longitudinal associations, clustering becomes a vital tool for determining reasonable regional divisions in stroke prevention. Nevertheless, this remains an unsolved issue that requires further investigation.
@@
In this work, we propose a novel method for clustering longitudinal associations between SDOH and stroke mortality. Generally, the clustering is implemented based on the similarity among SDOH-stroke mortality associations across counties in the U.S., and the resulting clustering outcomes can inform region-specific prevention measures for controlling stroke mortality. To achieve this, we utilize county-level longitudinal data on SDOH and stroke mortality in the U.S., treating the longitudinally observed data in each county as functional data (Ramsay and Silverman (2005); Wang et al. (2016)). Consequently, the concurrent associations between SDOH and stroke mortality are inherently functional objects, which can be effectively modeled using functional regression models (Yao et al. (2005); Wang et al. (2008); Zhu et al. (2012); Morris (2015); Kong et al. (2016); Zhou et al. (2023)), where the coefficients are also functional, capturing the relationship between SDOH and stroke mortality over time. For the clustering process, we model these functional coefficients using a finite mixture model, leading to a finite mixture of functional regression models. This approach enables the clustering of U.S. counties into different regions based on the relationship between their SDOH and stroke mortality.
Related Literature
Our task essentially differs from the common clustering procedures for longitudinal data (Chiou and Li (2007); Peng and Müller (2008); McNicholas and Murphy (2010); Jacques and Preda (2013); Komárek and Komárková (2013); Jacques and Preda (2014); Tang and Qu (2016)), which primarily apply to observable longitudinal outcomes. In contrast, the county-level longitudinal associations between SDOH and stroke mortality in our study are unobservable. To cluster these associations, one might consider applying finite mixtures of regression models (Peel and MacLahlan (2000); Khalili and Chen (2007); Khalili and Lin (2013)) to the SDOH and stroke mortality longitudinal data. However, this approach does not account for longitudinal changes in the associations within the clustering procedures, which could lead to unreliable clustering outcomes due to the disregard of longitudinal signals.
@@
In light of functional data analysis (Ramsay and Silverman (2002, 2005); Wang et al. (2016)), some studies proposed a finite mixture of functional regression models (Yao et al. (2011); Lu and Song (2012); Wang et al. (2023)) to account for time-varying changes within mixture regression models. However, these methods do not address the issue of collinearity among covariates within their regression models. In our case, the collection of SDOH served as functional covariates, exhibits significant collinearity due to their derivation from multiple domains (e.g., social, economic, and environmental domains). Ignoring collinearity among the SDOH data may lead to misspecification in the functional regression model, potentially compromising the accuracy of the resultant clustering outcome.
@@
To enhance interpretability, existing studies analyzing SDOH-stroke mortality associations often adopt a preselection step on the SDOH covariates (Tsao et al. (2022, 2023)), as their number is usually large, containing hundreds of variables. Without a proper selection of SDOH covariates, direct clustering using the existing methods (Yao et al. (2011); Lu and Song (2012); Wang et al. (2023)) may be statistically inefficient due to the complex structure of longitudinal SDOH data, which confronts high-dimensional challenges and complex functional structures simultaneously. To solve this issue, one may perform variable selection in functional regression models (Wang et al. (2008); Goldsmith and Schwartz (2017); Ghosal et al. (2020); Aneiros et al. (2022); Cai et al. (2022)) prior to clustering. However, these methods typically assume that associations between covariates and responses are invariant across samples, an assumption that limits their applicability to cases with heterogeneous associations. Moreover, performing selections of SDOH prior to clustering may also introduce biases for the subsequent clustering outcome, as the selection of SDOH is unrelated to the clustering process. To overcome theses challenges, it would be beneficial to connect the variable selection of functional covariates to the clustering of longitudinal associations; yet this topic is rarely discussed in the literature.
Contributions
@@
In this article, we introduce a novel method for simultaneous clustering and variable selection of longitudinal associations between Social Determinants of Health (SDOH) and stroke mortality across U.S. counties. Our approach utilizes a Finite Mixture of Functional Linear Concurrent Models (FMFLCM), incorporating a new class of penalties that pursue both smoothness and sparsity. These penalties are designed to borrow information across different clusters within the functional regression models, enhancing statistical efficiency for both clustering and variable selection while addressing collinearity among functional covariates. For estimation, we develop a novel Regularized Expectation-Maximization (REM) algorithm that integrates the proposed penalties. This method facilitates the clustering of longitudinal associations while embedding a variable selection step for functional covariates. As a result, we can iteratively update the cluster memberships and the selected covariates within the REM algorithm, mitigating biases that may arise from selected covariates during the clustering process. Through this methodology, we provide a new data-driven approach for county-level regional division, aiming to deliver insights into region-specific stroke prevention measures for the U.S. counties.
Organization
@@
The remainder of this article is organized as follows: In Section 2 we introduce the SDOH and stroke mortality dataset and then proceed to describe some of their data features in Section 2.1. Subsequently, we present the Finite Mixture of Functional Linear Concurrent Models (FMFLCM) in Section 2.2, followed by a demonstration of the sparsity- and smoothness-pursued penalties in Section 2.3, as well as the Regularized Expectation-Maximization (REM) algorithm in Section 2.4. In Section 3 we conduct simulation studies to compare the proposed methods with competing approaches in terms of clustering performance, variable selection, and parameter estimation. In Section 4 we apply the proposed clustering method to our dataset and present the clustering results, alongside the variable selection for the SDOH covariates. We conclude with a discussion in Section 5. The codes and datasets are publicly available at https://github.com/fl81224/Functional-Clustering-of-Longitudinal-Associations.
@@ 2.1 Data Source
Figure 1: The correlation matrix for the SDOH variables (indexed by numbers). The correlation of the longitudinal data between the $j$th SDOH and $j^{\prime}$th SDOH is calculated as $C_{j,j^{\prime}}/\sqrt{C_{j,j}C_{j^{\prime},j^{\prime}}}$, where $C_{j,j^{\prime}}=n^{-1}\sum_{i=1}^{n}S_{i}^{-1}\sum_{s=1}^{S_{i}}\big{\{}X_{ij}(t_{is})-\hat{\mu}_{j}(t_{is})\big{\}}\big{\{}X_{ij^{\prime}}(t_{is})-\hat{\mu}_{j^{\prime}}(t_{is})\big{\}}$, with $\hat{\mu}_{j}(\cdot)$ being the estimated mean function for the $j$th county. Here, $X_{ij}(t_{is})$ is the observed data of the $j$th SDOH at the $i$th county on year $t_{is}$, for $i=1,\dots,n$, $j=1,\dots,p$, and $s=1,\dots,S_{i}$. The definition of $X_{ij}(t_{is})$, $n$, $p$, and $S_{i}$ can be found in Section 2.2.
In 2020, the Agency for Healthcare Research and Quality (AHRQ) released a SDOH database to enhance understanding of community-level factors, healthcare quality and delivery, and individual health. The SDOH database contains yearly records of $345$ SDOH collected from 3226 counties from 2009 to 2018. These SDOH are classified into five domains: (1) social context, such as age, race and ethnicity, and veteran status, (2) economic context, such as income and unemployment, (3) education, (4) physical infrastructure, such as housing, food insecurity, and transportation, and (5) health care contexts, such as health insurance coverage and health care access. To study the association between SDOH and stroke mortality, we connect the SDOH database with a stroke mortality data provided by the Interactive Atlas of Heart Disease and Stroke at the CDC on the county level. The stroke mortality database was originally compiled from two data sources: (1) the National Vital Statistics System at the National Center for Health Statistics and (2) the hospital discharge data from the Centers for Medicare & Medicaid Services’ Medicare Provider Analysis and Review (MEDPAR) file. All data and materials used in this analysis are publicly available at the AHRQ website: https://www.ahrq.gov/sdoh/index.html and CDC website https://www.cdc.gov/dhdsp/maps/atlas/index.htm. Since the AHRQ database is HIPAA (Health Insurance Portability and Accountability Act) compliant, our data do not require to be reviewed by an institutional review board.
It is worth noting that the SDOH dataset contains missing values. Following the approach of a previous study on the dataset (Son et al. (2023)), we exclude the SDOH variables that have a missing proportion of more than 60%. The average missing proportion of the remaining variables is 3.6%. For these missing values, we employ a K-Nearest Neighbors (KNN) method (Kowarik and Templ (2016)) to impute the SDOH data, ensuring that the SDOH and stroke mortality data are aligned in time for each county. The detailed implementation of the KNN is provided in Part A of the Supplementary Material.
@@
Figure 2: Statewise and rural-urban Pearson correlations between the stroke mortality and PAPL. Panels (A) and (C) are the geographical maps of correlations for the years 2009, 2013, and 2018, respectively, for two types of division strategies. Panels (B) and (D) are the statewise and rural-urban longitudinal correlations from 2009–2018, respectively.
As mentioned in the Introduction, significant collinearity may exist among the SDOH data. This phenomenon is observed in our dataset, where the longitudinal data between some of the SDOH exhibit significant correlations, as illustrated in Figure 1. Besides, coinciding with the previous studies (Villablanca et al. (2016); He et al. (2021); Son et al. (2023)), evident regional disparities and longitudinal variations can also be observed in the SDOH-stroke mortality associations from our dataset. To demonstrate this, we calculate the yearly Pearson correlation between stroke mortality and the percentage of Asian and Pacific Island language speakers (PAPL), an important sociocultural factor in the SDOH dataset. To calculate the yearly correlations, we adopt difference strategies to divide counties in the U.S. into several regions, followed by calculating the correlations in each divided region. Specifically, we illustrate the correlations based on two types of divisions: statewise division (Zelko et al. (2023)) and rural-urban division (Son et al. (2023)). The yearly correlations from 2009–2018 for each state, or for rural and urban areas, are presented in panels (B) and (D) of Figure 2, respectively. Besides, we also present the geographical maps of correlations for the years 2009, 2013, and 2018 in panels (A) and (C) of Figure 2, respectively, for two types of division strategies.
@@
From the maps in Figure 2, we observe significant regional disparities in the correlations among different states as well as between rural and urban areas. These correlations all exhibit changes over time. In addition, we find that the longitudinal correlations in rural and urban areas are quite different from the correlations from the states, as displayed in panels (B) and (D) of Figure 2. These results suggest that the outcomes of longitudinal associations are sensitive to the region division strategy, and different region divisions may lead to inconsistent conclusions for the stroke management. To address this issue, we require a reasonable data-driven method for the region division in exploring SDOH-stroke mortality associations.